A child goes down a slide, starting from rest. If the length of the slide is 2 m and it takes the child 3 seconds to go down the slide, what is the child’s acceleration?
1
Expert's answer
2020-09-28T08:09:18-0400

We can find the child's acceleration from the kinematic equation:


"d = v_0t+ \dfrac{1}{2}at^2,"

here, "d = 2 \ m" is the distance traveled by the child, "v_0 = 0 \ ms^{-1}" is the initial velocity of the child, "a" is the child's acceleration and "t = 3 \ s" is the time during which the child reaches down the slide.

Then, from this equation we can calculate the child's acceleration:


"a = \dfrac{2d}{t^2} = \dfrac{2 \cdot 2 \ m}{(3 \ s)^2} = 0.44 \ ms^{-2}."

Answer:

"a = 0.44 \ ms^{-2}."